OpenStudy (anonymous):

integration by parts integral of x(x+1)^5 dx

OpenStudy (anonymous):

\[\int\limits x(x+1)^5\]

OpenStudy (kainui):

I think you might be better off by using a substitution, such as x+1=u.

OpenStudy (vishweshshrimali5):

@givanna618 , @Kainui is perfectly correct

OpenStudy (vishweshshrimali5):

Though even substituting u = x will work

OpenStudy (vishweshshrimali5):

But why bother using u then ?

OpenStudy (tkhunny):

You can always expand it. \((x+1)^{5} = x^{5} + 5x^{4} + 10x^{3} + 10x^{2} + 5x + 1\)

OpenStudy (kainui):

Yeah but that's really not necessary...

OpenStudy (vishweshshrimali5):

\[\large{\int x(x+1)^5 \ dx}\] \[\large{= \int (u-1)u^5 du}\] \[\large{= \int u^6 du - \int u^5 du}\] That is why @kainui suggested what he suggested

OpenStudy (vishweshshrimali5):

Using, u = x+1, ensures that you don't have to use binomial theorem :)

OpenStudy (vishweshshrimali5):

It gets easier that way

OpenStudy (vishweshshrimali5):

Great work @Kainui :D

OpenStudy (tkhunny):

Necessary? It's better than a brain cramp. It's absolutely the hardest way to do it, but it is a way to do it. If the alternative is to punt, why not do it a hard way?

OpenStudy (vishweshshrimali5):

Very wise words @tkhunny :)

OpenStudy (vishweshshrimali5):

Well @givanna618 now you have 2 different methods of solving the question thanks to @tkhunny and @Kainui . You can use whichever you like

OpenStudy (vishweshshrimali5):

And you can also use integration by parts

OpenStudy (vishweshshrimali5):

Though I would not prefer it ^^^

OpenStudy (tkhunny):

"by parts" was my third option.

OpenStudy (vishweshshrimali5):

:) By parts is a very general method. It can be used in many many integrals.

OpenStudy (vishweshshrimali5):

Sometimes I really wish there was an option of giving more than 1 medal.